A small but measurable current of 8.9 × 10-10 A exists in a copper wire whose diameter is 2.2 mm. The number of charge carriers per unit volume is 8.49 × 1028 m-3. Assuming the current is uniform, calculate the (a) current density and (b) electron drift speed.
Answer #1

Solution)

Given,

Current, i=8.9*10^-10 A

Radius r=1.1*10^-3 A

n=8.49*10^28 m^-3

Part a)

We know,

Current density, J=i/pi*(r)^2=(8.9*10^-10)/(pi*0.0011^2)

J=7.35*10^-6 A/m^2

=======

Part b)

Drift speed, V=J/ne=(7.35*10^-6)/(8.49*10^28)*(1.6*10^-19)

V=5.41*10^-16 m/s

=======